OpenStudy (anonymous):

Write an equation for the line that is parallel to y-5x+2=0. The line you are looking for passes through (-3, -8). Write the line in standard form.

OpenStudy (anonymous):

Ax +By= C this is standard form like the last problem slope=5 so what would you do? look at last problem? tell me what you think

OpenStudy (anonymous):

so the line is 5x-y=-2

OpenStudy (anonymous):

but how do the -3 and -8 factor into this problem?

OpenStudy (anonymous):

5x-y=2 would be the y-5x+2=0, but thats not the one your doing. What did we do with the (-5,1) from last problem